OpenStudy (anonymous):

x2-2x +4 for x=5 when x=5,x2-2x+4=

OpenStudy (diyadiya):

Substitute x=5 into the equation x2-2x +4

OpenStudy (diyadiya):

\[(5)^2-2(5)+4\]

OpenStudy (anonymous):

\[\Large \begin{array}{l} {x^2} - 2x + 4\\ x = 5\\ = {5^2} - 2 \cdot 5 + 4\\ = 25 - 10 + 4\\ = ? \end{array}\]

OpenStudy (anonymous):

Since you're new to algebra Dee2012, a simple polynomial such as the above is evaluated by simply replacing the co-efficient with its value. So everywhere you see x, replace it with 5, and evaluate the expression.